8(b) A trading company buys and sells 10,000 bottles of pain-balm every year. The

company’s cost of placing an order of pain-balm is Rs. 100. The holding cost per

bottle on inventory is Rs. 0.3. (5)

(i) Determine the optimum order quantity and inventory cycle time for the pain-balm,

bottles.

(ii) How many orders should be placed each year?


1
Expert's answer
2021-05-07T09:40:28-0400

i) The optimum order quantity kk is the ratio between the cost of placing an order mm and the holding cost "p"


"k=\frac mp=\frac{100}{0.3}=333 bottles"


inventory cycle time tt is the ratio between the optimum order quantity kk and the quantity of the bottles "n"


"t=\frac kn=\frac{333}{10000}=\frac{1}{30}" year (or 1212 days).


ii) Number of orders ss is the inverse of the inventory cycle time t,t,


"s=\frac 1t=30" orders per year.